OpenStudy (anonymous):

find integral {e^sin(2x)cos(2x) dx

OpenStudy (anonymous):

lols you typed in the wrong section

OpenStudy (anonymous):

^

OpenStudy (anonymous):

sub u=sin(2x) lols eventually u will need to learn to do them yourself :P

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4} \]

OpenStudy (anonymous):

lol elec, you're starting off right, proceed :)

OpenStudy (anonymous):

remember to change limits to the new variable

OpenStudy (anonymous):

not sure but i got 1/2 e^sin(2x)

OpenStudy (anonymous):

well easiet way to check if you did the integral right is to differentiate it

OpenStudy (anonymous):

looks correct

OpenStudy (anonymous):

;-) im getting it finally

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

excellent ^_^